An object is launched straight up into the air at an initial velocity of 64 feet per second. It is launched from a height of 6 feet off the ground. Its height H, in feet, at t seconds is given by the equation H= -16tsquared + 64t + 6. Find all the times t that the object is at a height of 54 feet off the ground.
Answer by stanbon(75887)

An object is launched straight up into the air at an initial velocity of 64 feet per second. It is launched from a height of 6 feet off the ground. Its height H, in feet, at t seconds is given by the equation
H= -16tsquared + 64t + 6.
Find all the times t that the object is at a height of 54 feet off the ground.
-----
Solve: -16t^2+64t+6 = 54
-16t^2+64t-48 = 0
-16(t^2-4t+3) = 0
t^2-4t+3 = 0
(t-3)(t-1) = 0
time = 1 second to get to 54 feet on the way up.
and t = 3 seconds to be at 54 feet on the way down.
====================================
Cheers,
Stan H.